OpenStudy (anonymous):

How can √3 cos a−sin a=2((√3/2) cos a− (1/2) sin a) become 2(cos θ cos a - sin θ sin a) and then become 2 cos (a + θ) Can someone explain this please?

OpenStudy (asnaseer):

do you know how to expand out \(\cos(a+\theta)\)?

OpenStudy (anonymous):

Ah that's where it comes from ok, sorry I didn't see that. But what happens to the \[\sqrt{3}/2 \] and the 1/2 ?

OpenStudy (asnaseer):

expand \(\cos(a+\theta)\) and then see which terms equate to \(\frac{\sqrt{3}}{2}\) and \(\frac{1}{2}\) - you should then be able to spot what to do next

OpenStudy (asnaseer):

i.e. you have:\[2(\frac{\sqrt{3}}{2}\cos(a)-\frac{1}{2}\sin(a))=2(\cos(\theta)\cos(a)-\sin(\theta)\sin(a))\]look at this identity and see what it tells you

OpenStudy (anonymous):

ok... I think I've got it

OpenStudy (asnaseer):

great! :)

OpenStudy (anonymous):

Thanks

OpenStudy (asnaseer):

yw :)